The total pressure is 710 torr what are the partial pressures if the gas present consists of 1.2 mol of N2, 0.26 mol of O2 and 0.068 mol of Ar.

Partial pressure of N2= ???? torr

Partial pressure O2= ???? torr

Partial pressure Ar= ??? torr
Answer #1

Solution:

According to Rault:

Partial pressure (pi) is calculated from total pressure (P) as:

pi = xi P

Where, xi= mole fraction

Total number of moles (n) of N2, O2 and Ar

n = 1.2 + 0.26 + 0.068 =1.528

Mole fraction = number of mole / total number of moles

A) Mol fraction of N2 = 1.2 / 1.528

Then, partial pressure = mole fraction x total pressure

p (N2) = (1.2 / 1.528) x 710 torr = 557.59 torr

B) Mole fraction of O2 = 0.26 / 1.528

p (O2) = (0.26/1.528) x 710 torr =120.81 torr

C) Mole fraction of Ar = 0.068 / 1.528

p(Ar) =(0.068/1.528) x 710 torr = 31.60 torr